OpenStudy (anonymous):

What is the derivative of x+(4/x) using the limit process?

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{[x+h+\frac{4}{x+h}]-[x+\frac{4}{x}]}{h} \\ =\lim_{h \rightarrow 0}\frac{1}{h}[x+h+\frac{4}{x+h}-x-\frac{4}{x}]\] We want to try to cancel that lonely h on the bottom (the one outside the brackets)

OpenStudy (freckles):

first of all do you see anything that cancels inside the [ ]

OpenStudy (anonymous):

The x cancels out

OpenStudy (anonymous):

what i ended up with was \[(x(x^2+xh+h^2+4)-(x+h)(x^2+4)) / (x+h)(x)\]

OpenStudy (anonymous):

but so far ive gotten stuck

OpenStudy (freckles):

so you tried to combine the fractions inside the brackets after x-x=0

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[\frac{1}{h}[\frac{h(x)(x+h)+4(x)-4(x+h)}{x(x+h)}]\]

OpenStudy (freckles):

so on top there we have hx(x+h)+4x-4x-4h

OpenStudy (freckles):

you should see something else that cancels

OpenStudy (anonymous):

the 4xs, but this is just the continuation of crossing out the original Xs right? cause I think I just did it in a way that made the problem a whole lot more complicated for me

OpenStudy (freckles):

Yeah the 4x's cancel I can't really tell how you got your fraction up there... Do you understand how I got mine fraction?

OpenStudy (anonymous):

No I'm not really sure... I think I have an idea

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{[x+h+\frac{4}{x+h}]-[x+\frac{4}{x}]}{h} \\ =\lim_{h \rightarrow 0}\frac{1}{h}[x+h+\frac{4}{x+h}-x-\frac{4}{x}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[\cancel{x} +h+\frac{4}{x+h}-\cancel{x}-\frac{4}{x}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[h+\frac{4}{x+h}-\frac{4}{x}] \\ \\ \text{ lcm of the bottoms is x(x+h) } \\ =\lim_{h \rightarrow 0}\frac{1}{h}[h \frac{x(x+h)}{x(x+h)}+\frac{4}{x+h}\frac{x}{x}-\frac{4}{x} \frac{x+h}{x+h}] \\ \lim_{h \rightarrow 0}\frac{1}{h}[ \frac{hx(x+h)+4x-4(x+h)}{x(x+h)}]\]

OpenStudy (freckles):

That is how I combined the fractions inside the []

OpenStudy (freckles):

now after canceling those 4x's you will that all that remains are terms in the numerator that have a common factor h so you can factor that h on top and cancel it with the bottom lonely bottom h outside the [ ]

OpenStudy (anonymous):

Okay I just figured it out at the same time you sent me that!

OpenStudy (freckles):

one you get rid of that factor h there that is on bottom (with the top factor h) you will be able to replace remaining h's with 0's and find f'(x)

OpenStudy (anonymous):

Wait I'm sorry Im confused so I have \[hx(x+h)-4h / h\] and I am factoring out ...?

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{[x+h+\frac{4}{x+h}]-[x+\frac{4}{x}]}{h} \\ =\lim_{h \rightarrow 0}\frac{1}{h}[x+h+\frac{4}{x+h}-x-\frac{4}{x}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[\cancel{x} +h+\frac{4}{x+h}-\cancel{x}-\frac{4}{x}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[h+\frac{4}{x+h}-\frac{4}{x}] \\ \\ \text{ lcm of the bottoms is x(x+h) } \\ =\lim_{h \rightarrow 0}\frac{1}{h}[h \frac{x(x+h)}{x(x+h)}+\frac{4}{x+h}\frac{x}{x}-\frac{4}{x} \frac{x+h}{x+h}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[ \frac{hx(x+h)+4x-4(x+h)}{x(x+h)}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[\frac{xh(x+h)-4h}{x(x+h)}] \\ =\lim_{h \rightarrow 0}\frac{h}{h} [\frac{x(x+h)-4}{x(x+h)}]\]

OpenStudy (freckles):

now remember h/h=1 so now you have gotten rid of the thingy that would make the bottom 0 so replace remaining h's with 0

OpenStudy (anonymous):

Okay so I'm just left with \[\left( x^2-4\right)/x^2\]

OpenStudy (freckles):

looks good

OpenStudy (anonymous):

why wouldn't I just cancel out the x(x+h)

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{[x+h+\frac{4}{x+h}]-[x+\frac{4}{x}]}{h} \\ =\lim_{h \rightarrow 0}\frac{1}{h}[x+h+\frac{4}{x+h}-x-\frac{4}{x}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[\cancel{x} +h+\frac{4}{x+h}-\cancel{x}-\frac{4}{x}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[h+\frac{4}{x+h}-\frac{4}{x}] \\ \\ \text{ lcm of the bottoms is x(x+h) } \\ =\lim_{h \rightarrow 0}\frac{1}{h}[h \frac{x(x+h)}{x(x+h)}+\frac{4}{x+h}\frac{x}{x}-\frac{4}{x} \frac{x+h}{x+h}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[ \frac{hx(x+h)+4x-4(x+h)}{x(x+h)}] \\ =\lim_{h \rightarrow 0}\frac{1}{h}[\frac{xh(x+h)-4h}{x(x+h)}] \\ =\lim_{h \rightarrow 0}\frac{h}{h} [\frac{x(x+h)-4}{x(x+h)}] \\ \text{ you could write this as } \\ =\lim_{h \rightarrow 0}\frac{h}{h} [1-\frac{4}{x(x+h)}]\]

OpenStudy (freckles):

That will still lead to the same answer...

OpenStudy (anonymous):

ooooh okay thankss!

OpenStudy (freckles):

\[\frac{x^2-4}{x^2}=\frac{x^2}{x^2}-\frac{4}{x^2}=1-\frac{4}{x^2}\]

OpenStudy (anonymous):

hahaha Idk why I didn't see that earlier

OpenStudy (anonymous):

thank you so much, this was a tough one